OpenStudy (anonymous):

Help with a system of equations? a + s = 150 45a + 28.5 = 4770

OpenStudy (anonymous):

it it 5 or s in the first equation

OpenStudy (anonymous):

s

OpenStudy (anonymous):

ok from the second equation can you find the value of a ?????

OpenStudy (anonymous):

I'm trying to figure out how... divide?

OpenStudy (anonymous):

use calculator

OpenStudy (anonymous):

ok let me make it easy for you

OpenStudy (anonymous):

but I have to use substitution or elimination

OpenStudy (anonymous):

substitution

OpenStudy (anonymous):

as you find the value of " a " from the second equation put that value in the first equation and you will get " s " thats it

OpenStudy (anonymous):

ohhh okay. I understand, thank you!

OpenStudy (anonymous):

welcome

OpenStudy (phi):

45a + 28.5 = 4770 did you add -28.5 to both sides, like this: 45 a + 28.5 + -28.5 = 4770 + -28.5 now simplify. what do you get ?

OpenStudy (anonymous):

45a = 4741.5?

OpenStudy (anonymous):

now divide both sides by 45 what you get???

OpenStudy (phi):

yes. now do as rizwan suggests: divide both sides by 45 45 a/45 = 4741.5/45 on the left side 45/45 is 1 so you get 1*a or just a

OpenStudy (anonymous):

thank you so much! I appriciate your help!